OpenStudy (anonymous):

A theater has 20 rows of chairs. The first row has 16 chairs, the second has 18, the third has 20, and so on. In total, the theater seats _____ people. ****So I think the pattern is add 2 for each row.... But idk how to solve it... is there like an equation or something???

OpenStudy (anonymous):

Hi!! Thanks for coming to help!! :) Do you understand this? ;)

jimthompson5910 (jim_thompson5910):

You have the sequence: 16, 18, 20, 22, 24, ... You want to add up the terms: 16+18+20+22+24+...(Twentieth Term) You can use the formula Sn = n*(a1 + an)/2 where n is the number of terms, a1 is the first term and an is the nth term

OpenStudy (anonymous):

ohhhh, so i forgot though, what is the n mean?? is that 20?? since there are 20 rows?

jimthompson5910 (jim_thompson5910):

n = 20, yes

jimthompson5910 (jim_thompson5910):

Also remember that an = a1+ (n-1)*d where d is the common difference

OpenStudy (anonymous):

so a20=(16)+(20-1)(2) is that right so far?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

so i get 54??

jimthompson5910 (jim_thompson5910):

now plug this into the formula Sn = n*(a1 + an)/2 S20 = 20*(16 + 54)/2 ...

OpenStudy (anonymous):

so this was the first formula you listed right?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

And i get (20)(70)/2 which equals 1400/2 which = 700 ???

jimthompson5910 (jim_thompson5910):

you nailed it

jimthompson5910 (jim_thompson5910):

If you want to do it the long way, then list out the first 20 terms: 16, 18, 20, 22, 24, 26, 28, 30, 32, 34, 36, 38, 40, 42, 44, 46, 48, 50, 52, 54 Then add them up 16+18+20+22+24+26+28+30+32+34+36+38+40+42+44+46+48+50+52+54 = 700 So we get the same answer

OpenStudy (anonymous):

yay!!! :) Thanks so much @jim_thompson5910 !!! you rock!! :) oh sooo i can do it both ways and i get the same answer?

jimthompson5910 (jim_thompson5910):

either way works, the formula is the best and easiest way, but it's good to know the formula works (by confirming it with the long way)

OpenStudy (anonymous):

oh okay!! That's good to know :) Thank you!! :)

jimthompson5910 (jim_thompson5910):

yw